Q3. Determine its average unit strain in a reinforced concrete column that is 40ft tall, and under lad shortens 1/2 in.

Q4. Determine the average unit strain in a wood post 25ft tall that shortens 1/4in under load.

Q5. A 200ft long steel cable is loaded in tension until the average unit strain is 0.004. Determine the total elongation under this load.

Q6. in explaining the Behavior of concrete under load, one popular theory indicates that concrete will crush when the compressive unit strain reaches 0.003. If a 12in tall concrete test cylinder is failed in a testing machine, what will be its height when it fails?
Answer #1

3) Initial height of column = 40 ft = 480 in

Shortening length = 0.5 in

Compressive average unit strain = 0.5/480 = 0.001042

4) Initial Height of post = 25ft = 300 in

Shortening length = 0.25 in

Compressive average unit strain = 0.25/300 = 0.000833

5) Initial length of cable = 200 ft = 2400 in

Average unit strain = 0.004

Total elongation = 0.004 x 2400 = 9.6 in

(6) Initial height of cylinder = 12 in

Compressive unit strain = 0.003

Shortening height = 0.003 x 12 = 0.036 in

Final height at failure = 12 - 0.036 = 11.964 in